one number is eight less than six times a second number. the sum of the two numbers is thirty-four. find each number.
I tried to write this problem out like
x-8(6+x)=34 but i am pretty sure that is not right can you email me back the answer with details on how to lay this problem out and solve it thank you
Jennifer

Answer by josgarithmetic(39838)

"one number", x;
"a second number", y.
x=-8+6y.
x+y=34.

Looking for a way to proceed a bit further, substitute x statement into the second equation: (-8+6y)+y=34. The rest will flow smoothly.